# Quantum Normalizing Flows for Anomaly Detection

### v2

@@ II.2 Normalizing Flows
Figure 1: Visualization of a normalizing flow $f$. It is a bijective transformation $f$ of an arbitrary distribution to e.g. a normal distribution.
@@ II.3.1 Quantum anomaly detection
One of the first approaches to formulate anomaly detection on quantum computers has been proposed in PhysRevA.97.042315 . The authors mainly build up on a kernel PCA and a one-class support vector machine, one of the approaches we will introduce in more detail later. So-called change point detection has been analyzed in PhysRevLett.119.140506 ; PhysRevA.98.052305 ; fanizza2023ultimate . In PhysRevA.99.052310 , Liang et.al. propose anomaly detection using density estimation. Therefore it is assumed that the data follows a specific type of distribution, in its simplest form a Gaussian mixture model. In GUO2023129018 the Local Outlier Factor algorithm (LOF algorithm) LOF2000 has been used and remapped to a quantum formulation. As explained later, the LOF algorithm contains three steps, (a) determine the k-distance neighborhood for each data point $x$, (b) compute the local reachability density of $x$, and (c) calculate the local outlier factor of $x$ to judge whether $x$ is abnormal. In GUO2022127936 the authors propose an efficient quantum anomaly detection algorithm based on density estimation which is driven from amplitude estimation. They show that their algorithm achieves exponential speed up on the number of training data points $M$ over its classical counterpart. Besides fundamental theoretical concepts, many works do not show any experiments on real datasets and are therefore often limited to very simple and artificial examples. Also the generated quantum gate sequences can require a large amount of qubits and they lead to reasonable large code lengths which is suboptimal for real-world scenarios.
In the following we will summarize three classical and well established methods which we will later use for a direct comparison to our proposed Quantum-Flow.
@@ II.3.2 Isolation forests
An isolation forest is an algorithm for anomaly detection which has been initially proposed by Liu et.al. liu2008isolation . It detects anomalies using characteristics of anomalies, i.e. being few and different. The idea behind the isolation forest algorithm is that anomalous data points are easier to separate from the rest of the data. In order to isolate a data point, the algorithm generates partitions on the samples by randomly selecting an attribute and then randomly selecting a split value in a valid parameter range. The recursive partitioning leads to a tree structure and the required number of partitions to isolate a point corresponds to the length of the path in the tree. Repeating this strategy leads to an isolation forest and finally, all path lengths in the forest are used to determine an anomaly score. The isolation forest algorithm computes the anomaly score s(x) of an observation x by normalizing the path length h(x):
@@ III.1 Optimization
Optimization is based on quantum architecture search. The name is borrowed and adapted from Neural Architecture Search (NAS) Miikkulainen2020 ; xie2018snas , which is devoted to the study and hyperparameter tuning of neural networks. Many QAS-variants are focussed on discrete optimization and exploit optimization strategies for non-differentiable optimization criteria. In the past, variants of Gibbs sampling PhysRevResearchLi20 , evolutional approaches 9870269 , genetic algorithms RasconiOddi2019 ; PhysRevLett116.230504 and neural-network based predictors Zhang2021 have been suggested. A recent survey on QAS can be found in Zhu23ICACS . For this work, we rely on the former work RosOsb2023a proposing Monte Carlo Graph Search. The optimized loss function is in our case the Kullback-Leibler divergence, $D_{KL}(P\parallel Q)$ (see Equation (2)) which is the standard loss for many anomaly detection frameworks. We will also perform experiments using the cosine-similarity measure. As discussed later, this measure can be evaluated directly from quantum states via a SWAP test. Thus, our approach can be executed on a quantum computer with a sequence comprising of state preparation, the optimized quantum gates sequences and followed by evaluating the quantum cosine dissimilarity as proposed in 9605277 . Quantum Architecture search requires at the end (a) a pool of quantum gates to sample from, (b) a loss function and (c) some simple hyperparameters, e.g. the maximum length of the quantum gate sequence or a stopping criteria. Then different gate orders are sampled and evaluated. The used quantum architecture search algorithm is based on Monte Carlo graph search (MCGS) 8632344 ; Czech_Korus_Kersting_2021 and measures of importance sampling. MCGS is very efficient, since many operators for quantum computing act locally (e.g. the $H$-Gates, $X$-Gates, $Z$-Gates and more). Therefore many combinations of sampled unitary matrices representing the gate order are algebraically commutative. Thus, sampled gate orders can be summarized in a graph which has many cycles. Searching on such a graph for optimal codes allows to reduce the amount of unnecessary and redundantly sampled configurations, leading to an efficient sampling strategy. We refer to RosOsb2023a for further details.
@@ III.2 Evaluation Metrics
Figure 3: Optimized Normalizing Flow from an input distribution (magenta) to a target distribution (blue) and the comparison to a discrete Gaussian distribution.
@@ III.3 Quantum Flow Anomaly Detection
Figure 5: Obtained quantum gate sequences after optimization. Despite the competitive performance the resulting codes are reasonable small and efficient. (best viewed in color and zoom in)
@@
Figure 6: A Normalizing Flow has been trained on the first class for the iris and wine dataset, respectively. The crosses denote the TSNE-plot of the training data. The generated examples from sampling the normal distribution are shown in circles.
The quantum architecture search algorithm requires as input a pool of optional quantum gates to select from. In our experiments we used the Pauli-($X$, $Y$, $Z$) operations, as well as Hadamard-, cnot-, swap-, phase-shift-, and toffoli-gates. The phase shift gates have a continuous parameter $\theta$ we sampled with $\pi,\frac{\pi}{2},\frac{\pi}{4}$. Thus, only the discrete ordering and selection of quantum gates is optimized by the quantum architecture search. Table 1 summarizes the used datasets, the settings and the train and test splits. Figure 4 summarizes the resulting ROC-Curves and the area-under the ROC-Curve as a final quality measure. Additionally, we provide the ROC-Curves and the obtained results of isolation forests, the local outlier factor (LOF) and the single-class SVM. The final performance is also summarized in table 2. It is apparent that our proposed optimized Quantum Flows can achieve competitive performance. Additionally, the code is available as short and highly performant quantum gate sequence. The obtained quantum gate sequences are provided in Figure 5. In the following section we will discuss how the decision algorithm can be implemented on a quantum computer.
@@ III.5 Normalizing Flow as a generator
As already mentioned in Section 2.2, a Normalizing Flow can be used for several purposes, e.g. anomaly detection as shown in the previous part. Another application is to use the Normalizing Flow as a generator by sampling from the normal distribution and inverting the forward transformation. Thus, in the final experiment, a Normalizing Flow has been trained on the first class of both, the iris and wine dataset, respectively. Afterwards, samples from the normal distribution are generated and transformed using the inverse flow which leads to new samples in the original data. Figure 6 visualizes the tsne-plots for these datasets (in crosses) as well as generated examples (in circles) after sampling complex values from the normal distribution and inverting the learned forward mapping $U$. Note, that the backward transformation can lead to non-useful samples since only the distribution on the absolute values is optimized in the forward flow. Thus, sampling complex numbers can lead to unlikely examples. We therefore verify the samples by backprojecting them onto the normal distribution again and only select samples which have a small KL divergence. As expected, the generated samples are in the domain of the training data and generalize examples among them. Figure 6 shows in crosses the TSNE-data for the three classes and in circles the generated samples which are located around the first class label.
@@ IV Summary
In this work, quantum architecture search is used to compute a Normalizing Flow which can be summarized as a bijective mapping from an arbitrary distribution to a normal distribution. The optimization is based on the Kullback-Leibler-divergence. Once such a mapping has been optimized, it can be applied to anomaly detection by comparing the distribution of quantum measurements to the expected normal distribution. In the experiments we perform comparisons to three standard methods, namely isolation forests, the local outlier factor (LOF) and a single-class SVM. The optimized architectures show competitive performance despite being fully implementable on a quantum computer. Additionally we demonstrate how to use the Normalizing Flow as generator by sampling from a normal distribution and inverting the flow.

### v3

@@ II.2 Normalizing Flows
Figure 1: Visualization of a normalizing flow $f$. It is a bijective transformation $f$ of an arbitrary distribution (from a given dataset) to e.g. a normal distribution.
@@ II.3.1 Quantum anomaly detection
One of the first approaches to formulate anomaly detection on quantum computers has been proposed in PhysRevA.97.042315 . The authors mainly build up on a kernel PCA and a one-class support vector machine, one of the approaches we will introduce in more detail later. So-called change point detection has been analyzed in PhysRevLett.119.140506 ; PhysRevA.98.052305 ; fanizza2023ultimate . In PhysRevA.99.052310 , Liang et al. propose anomaly detection using density estimation. Therefore it is assumed that the data follows a specific type of distribution, in its simplest form a Gaussian mixture model. In GUO2023129018 the Local Outlier Factor algorithm (LOF algorithm) LOF2000 has been used and remapped to a quantum formulation. As explained later, the LOF algorithm contains three steps, (a) determine the k-distance neighborhood for each data point $x$, (b) compute the local reachability density of $x$, and (c) calculate the local outlier factor of $x$ to judge whether $x$ is abnormal. In GUO2022127936 the authors propose an efficient quantum anomaly detection algorithm based on density estimation which is driven from amplitude estimation. They show that their algorithm achieves exponential speed up on the number of training data points $M$ over its classical counterpart. Besides fundamental theoretical concepts, many works do not show any experiments on real datasets and are therefore often limited to very simple and artificial examples. Also the generated quantum gate sequences can require a large amount of qubits and they lead to reasonable large code lengths which is suboptimal for real-world scenarios GUO2023129018 ; PhysRevA.97.042315 .
In the following we will summarize three classical and well established methods which we will later use for a direct comparison to our proposed Quantum-Flow. For the experiments we used the implementation of these algorithms provided by matlab MATLAB . The optimization on the used datasets is very fast and takes less than a second on a standard notebook.
@@ II.3.2 Isolation forests
An isolation forest is an algorithm for anomaly detection which has been initially proposed by Liu et al. liu2008isolation . It detects anomalies using characteristics of anomalies, i.e. being few and different. The idea behind the isolation forest algorithm is that anomalous data points are easier to separate from the rest of the data. In order to isolate a data point, the algorithm generates partitions on the samples by randomly selecting an attribute and then randomly selecting a split value in a valid parameter range. The recursive partitioning leads to a tree structure and the required number of partitions to isolate a point corresponds to the length of the path in the tree. Repeating this strategy leads to an isolation forest and finally, all path lengths in the forest are used to determine an anomaly score. The isolation forest algorithm computes the anomaly score s(x) of an observation x by normalizing the path length h(x):
@@ III.1 Optimization
Optimization is based on quantum architecture search. The name is borrowed and adapted from Neural Architecture Search (NAS) Miikkulainen2020 ; xie2018snas , which is devoted to the study and hyperparameter tuning of neural networks. Many QAS-variants are focussed on discrete optimization and exploit optimization strategies for non-differentiable optimization criteria. In the past, variants of Gibbs sampling PhysRevResearchLi20 , evolutional approaches 9870269 , genetic algorithms RasconiOddi2019 ; PhysRevLett116.230504 and neural-network based predictors Zhang2021 have been suggested. A recent survey on QAS can be found in Zhu23ICACS . For this work, we rely on the former work RosOsb2023a proposing Monte Carlo Graph Search. The optimized loss function is in our case the Kullback-Leibler divergence, $D_{KL}(P\parallel Q)$ (see Equation (2)) which is the standard loss for many anomaly detection frameworks. We will also perform experiments using the cosine-similarity measure. As discussed later, this measure can be evaluated directly from quantum states via a SWAP test. Thus, our approach can be executed on a quantum computer with a sequence comprising of state preparation, the optimized quantum gates sequences and followed by evaluating the quantum cosine dissimilarity as proposed in 9605277 . Quantum Architecture search requires at the end (a) a pool of elementary quantum gates $\mathcal{OP}=\{O_{1},O_{2},\ldots\}$ to sample from, (b) a loss function and (c) some simple hyperparameters, e.g. the maximum length of the quantum gate sequence or a stopping criteria. Then different gate orders are sampled and evaluated. The used quantum architecture search algorithm is based on Monte Carlo graph search (MCGS) 8632344 ; Czech_Korus_Kersting_2021 and measures of importance sampling. MCGS is very efficient, since many operators for quantum computing act locally (e.g. the $H$-Gates, $X$-Gates, $Z$-Gates and more). Therefore many combinations of sampled unitary matrices representing the gate order are algebraically commutative. The MCGS-algorithms can be motivated from Monte-Carlo Tree Search (MCTS). It is a heuristic search algorithm for decision processes 8632344 . It makes use of random sampling and balances the exploration-exploitation dilemma in large search spaces. MCTS can be very efficient as it visits more interesting sub trees more often. Thus, it grows asymmetrically and focuses the search time on more relevant parts of the tree. For the MCGS, we start with the identity operator $\mathbb{I}$ and build quantum circuits by selecting elementary gates from a predefined set $\mathcal{OP}=\{O_{1},O_{2},\ldots\}$ of elementary quantum gates that we are allowed to apply. Due to the universality theorem nielsenQuantumComputationQuantum2000 it is possible to approximate any unitary matrix to arbitrarily good accuracy by using a sufficiently long product of such gates. The general idea is to grow a graphical model with nodes containing unitary matrices and edges encoding a unitary operator $O_{i}\in\mathcal{OP}$. The graph is initialized with the identity matrix $\mathbb{I}$ as root node. An operator $O_{i}$ is selected and applied to the root node. This yields a new node by multiplying the selected operator with the unitary matrix of the parent node (which is the identity matrix in the beginning). If the resulting unitary is already existing as node in the graph, a direct edge from the parent to the already existing node can be added. Else, a new node is generated and connected with the parent node. Thus, while growing the graph, the resulting unitary matrices are provided as graph nodes and the underlying quantum code can be computed by finding the shortest path from the root node to the target unitary and by collecting the operators along the edges of the path. Figure 5 is taken from RosOsb2023a and shows the general principle.
Figure 2: Tiny example graph of quantum circuits. The edges are labelled with elementary gates. The vertices are given by the unitary operator built by taking the product of the gates along the shortest path. (Image taken from RosOsb2023a )
Thus, each node is identified with a possible quantum circuit. Please note, that this graph contains cycles since identical quantum circuits have multiple representations with different gates and gate orders. The remaining challenge is to grow the graph in an efficient manner. Given a specific task, every node will receive a quality score which is used to compute a probability for its selection and development along this graph area. Poisson sampling is then exploited as the underlying sampling process for selecting a vertex to further develop. It is the basic paradigm of Monte Carlo Search doi:10.1080/01621459.1949.10483310 and adapted Gibbs sampling george1993variable to iteratively grow a graph. We refer to RosOsb2023a for further details.
@@ III.2 Evaluation Metrics
Figure 4: Optimized Normalizing Flow from an input distribution (magenta-cross) to a target distribution (blue-circle) and the comparison to a discrete Gaussian distribution (red-triangle).
@@ III.3 Quantum Flow Anomaly Detection
Iris1-2
[…]
Iris2-3
[…]
Wine1-2
[…]
Wine2-3
[…]
Figure 6: Obtained quantum gate sequences after optimization. Despite the competitive performance the resulting codes are reasonable small and efficient.
@@
Figure 7: A Normalizing Flow has been trained on the first class for the iris and wine dataset, respectively. The crosses denote the TSNE-plot of the training data. The generated examples from sampling the normal distribution are shown in circles and consistently fall into the learned domain.
The quantum architecture search algorithm requires as input a pool of optional quantum gates to select from. In our experiments we used the Pauli-($X$, $Y$, $Z$) operations, as well as Hadamard-, cnot-, swap-, phase-shift-, and toffoli-gates. The phase shift gates have a continuous parameter $\theta$ we sampled with $\pi,\frac{\pi}{2},\frac{\pi}{4}$. Thus, only the discrete ordering and selection of quantum gates is optimized by the quantum architecture search. Table 1 summarizes the used datasets, the settings and the train and test splits. Across all experiments, the optimization of the quantum circuits requires between 10 seconds and one minute. Please note, that this is only the (offline) training stage. Inference on a quantum computer does not require further optimization. Figure 5 summarizes the resulting ROC-Curves and the area-under the ROC-Curve as a final quality measure. Additionally, we provide the ROC-Curves and the obtained results of isolation forests, the local outlier factor (LOF) and the single-class SVM. The final performance is also summarized in table 2. It is apparent that our proposed optimized Quantum Flows can achieve competitive performance. Additionally, the code is available as short and highly performant quantum gate sequence. The obtained quantum gate sequences are provided in Figure 6. In the following section we will discuss how the decision algorithm can be implemented on a quantum computer.
@@ III.5 Normalizing Flow as a generator
As already mentioned in Section 2.2, a Normalizing Flow can be used for several purposes, e.g. anomaly detection as shown in the previous part. Another application is to use the Normalizing Flow as a generator by sampling from the normal distribution and inverting the forward transformation. Thus, in the final experiment, a Normalizing Flow has been trained on the first class of both, the iris and wine dataset, respectively. Afterwards, samples from the normal distribution are generated and transformed using the inverse flow which leads to new samples in the original data. Figure 7 visualizes the tsne-plots for these datasets (in crosses) as well as generated examples (in circles) after sampling complex values from the normal distribution and inverting the learned forward mapping $U$. Note, that the backward transformation can lead to non-useful samples since only the distribution on the absolute values is optimized in the forward flow. Thus, sampling complex numbers can lead to unlikely examples. We therefore verify the samples by backprojecting them onto the normal distribution again and only select samples which have a small KL divergence. As expected, the generated samples are in the domain of the training data and generalize examples among them. Figure 7 shows the TSNE-data for the three classes and in circles the generated samples which are located around the first class label.
@@ IV Summary
In this work, quantum architecture search is used to compute a Normalizing Flow which can be summarized as a bijective mapping from an arbitrary distribution to a normal distribution. The optimization is based on a Kullback-Leibler-divergence or the cosine dissimilarity. Once such a mapping has been optimized, it can be applied to anomaly detection by comparing the distribution of quantum measurements to the expected normal distribution. In the experiments we perform comparisons to three standard methods, namely isolation forests, the local outlier factor (LOF) and a single-class SVM. The optimized architectures show competitive performance despite being fully implementable on a quantum computer. Additionally we demonstrate how to use the Normalizing Flow as generator by sampling from a normal distribution and inverting the flow.